Integration by Parts Fractions


1.) ∫dy/(y^2+2y)


2.) ∫(x^3+x+2)/(x^2-1)dx


1
Expert's answer
2021-04-27T09:17:26-0400

1)

"\int\dfrac{dy}{y^2+2y}"

"\dfrac{1}{y^2+2y}=\dfrac{A}{y}+\dfrac{B}{y+2}=\dfrac{A(y+2)+By}{y(y+2)}"



"y=0: 2A=1=>A=\dfrac{1}{2}"

"y=-2: -2B=1=>B=-\dfrac{1}{2}"



"\int\dfrac{dy}{y^2+2y}=\dfrac{1}{2}\int\dfrac{dy}{y}-\dfrac{1}{2}\int\dfrac{dy}{y+2}"

"=\dfrac{1}{2}\ln|y|-\dfrac{1}{2}\ln|y-2|+C"

2)

"\int\dfrac{x^3+x+2}{x^2-1}dx=\int\dfrac{x^3-x+x+x+2}{x^2-1}dx"

"=\int xdx+\int\dfrac{2(x+1)}{(x+1)(x-1)}dx"

"=\dfrac{x^2}{2}+2\int\dfrac{dx}{x-1}=\dfrac{x^2}{2}+2\ln|x-1|+C"